The base of a triangle is 10 inches more than 2 times the height. If the area of the triangle is 14 square inches, find the base and height.
Answer by stanbon(75887)

The base of a triangle is 10 inches more than 2 times the height. If the area of the triangle is 14 square inches, find the base and height.
------
Let the height be "x":
Then the base is "2x+10"
---
Area = (1/2)(base)(height)
---
14 = (1/2)(2x+10)(x)
-----
2x^2 + 10x - 28= 0
---
x^2 + 5x - 14 = 9
(x+7)(x-2) = 9
Positive solution:
x = 2(height)
2x+10 = 14 (base)
===================
Cheers,
Stan H.